A body starts from rest, what is the ratio of the distance travelled by the body during the 4th and 3rd second?   


        (a)    7/5


(b)     5/7


(c)    7/3


(d)    3/7


Grade:10

1 Answers

Amit Saxena
10 years ago

(a)
    D_4/D_3 =   (0+a/2  (2×4-1))/(0+a/2  (2×3-1) )=7/5